OpenStudy (anonymous):

A rectangular prism has dimensions of 10 ft, 12 ft, and W ft. Its total surface area is 1076 ft2. What is W?

OpenStudy (anonymous):

woah prease delete those xD

OpenStudy (anonymous):

okay did it

OpenStudy (anonymous):

but okay lol so do you know the formula for SA for a rectangular prism?

OpenStudy (anonymous):

@cait1995

OpenStudy (anonymous):

nope

OpenStudy (anonymous):

2(wl+hl+hw) ???? @atreyu6s3x6

OpenStudy (anonymous):

A=2(wl+hl+hw)

OpenStudy (anonymous):

lmao dammit! yes haha so what do we plug in?

OpenStudy (anonymous):

1076= 2(w*10) + (12*10) +(12*w)

OpenStudy (anonymous):

but i don't know w..

OpenStudy (anonymous):

@atreyu6s3x6

OpenStudy (anonymous):

yes @cait1995 so then we multiply ech one of the parenthesis

OpenStudy (anonymous):

10w+120+12w @atreyu6s3x6

OpenStudy (anonymous):

ahh I'm running out of time

OpenStudy (anonymous):

yes so we add the like terms

OpenStudy (anonymous):

so 22w? @atreyu6s3x6

OpenStudy (anonymous):

1076=2(22w+120) distribute 2 1076=44w+240 subtract 240 836=44w divide by 44 w=19

OpenStudy (anonymous):

yes i did the rest since we were running low on time :)

OpenStudy (anonymous):

okay tahnks so much @atreyu6s3x6

OpenStudy (anonymous):

can you help with another? @atreyu6s3x6

OpenStudy (anonymous):

sure @cait1995